# The Effect of Sampling Temperature on Problem Solving in Large Language Models

### v2

@@ 1.1 Background
In addition, several inference hyperparameters can be adjusted to modify the LLM’s output at runtime. For example, hyperparameters like sampling temperature, top-k sampling, repetition penalty, and maximum token length all affect the LLMs output and performance [3, 4, 5].
Despite significant interest in LLMs and progress in LLM best practices, many open questions remain about optimal prompt-engineering techniques and inference hyperparameters for LLMs. To complicate matters, various local optima may exist for LLM, prompt type, and problem domain [5].
@@ 1.3 Choosing a Sampling Temperature
Practical guidelines for choosing a sampling temperature for a specific task or problem domain are often vague or anecdotal. Prompt engineering guides often provide hypothetical examples of optimal sample temperatures for various tasks. However, they rarely cite any sources or provide empirical evidence. [^1] A few empirical studies exist that indicate sampling temperature does have an effect on LLM performance on some types of problem-solving tasks (e.g., code generation, engineering exams, etc.) [11, 14, 12].
As a result, the current state of choosing the optimal sampling temperature for specific problems is largely based on guesswork, gut instinct, non-systematic experimentation, and iterative refinement. [^2] For example, OpenAI’s GPT-3.5 API allowed users to set the sampling temperature from 0.0 to 1.0 with a default of 0.7. GPT-4’s API expanded this range from 0.0 to 2.0 with a default of 1.0. No explanation from OpenAI has been provided for these default values or their change from GPT-3.5 to GPT-4 [15]. [^3] Even the GPT-4 Technical Report explains that the authors used their “best-guess” when choosing sampling temperatures while evaluating GPT-4 on various benchmarks. See Appendix A in the GPT-4 Technical Report [16].
@@ 2.1 Models
The LLM models used in this research project comprise nine widely-used foundational LLMs. To complement our analysis, we also conducted experiments using five prompts created using commonly used prompt-engineering techniques.
@@ 2.3 Process
First, we instructed GPT-3.5 to complete the 100-question (small) exam using the CoT prompt with temperatures ranging from 0.0 to 2.0 in increments of 0.1. This allowed us to determine the range of viable sampling temperatures to explore.
@@ 3 Results
Llama 2 7B did not perform better than statistically random guesses. Its poor performance was due to generating incorrectly formatted answers (39%) and correctly formatted but incorrect answers (36%). Its all-or-nothing behavior at a temperature of 0.0 versus more random behavior at 0.1 to 1.0 led to a much lower, yet still non-significant, p-value.
@@
Finally, we analyzed the performance of GPT-3.5 using the CoT prompt on all ten exams. Accuracy remained stable for all temperatures across all problem domains based on visual analysis. However, the LSAT-AR and SAT-Math exams showed statistically significant differences in the Kruskal-Wallis p-values. See Figure 5 and Table 5.
We performed the Dunn-Bonferroni test on the LSAT-AR and SAT-Math results [49]. It revealed that the all-or-nothing behavior of responses generated at a temperature of 0.0 versus the more random responses from 0.1 to 1.0 caused the anomaly. The correct-answer accuracy for each individual problem varied significantly when evaluated pair-wise across temperatures. However, the average accuracy for all problems remained similar across temperatures.
@@ 4.1 Interpretation
Based on these results, changes in temperature from 0.0 to 1.0 do not have a statistically significant effect on the problem-solving performance of LLMs. These results appear to generalize across LLMs, prompt-engineering techniques, and problem domains. However, there were a few potential exceptions to these general findings.
Therefore, we recommend setting an LLM’s sampling temperature to 0.0 for problem-solving tasks. This temperature maximizes reproducibility without compromising accuracy. In addition, it helps avoid the inevitable dropoff in performance that occurs beyond a temperature of 1.0. However, exceptions to this guidance should be taken into consideration.
@@ 4.2 Limitations
Fourth, we had to limit the sampling temperature we explored from 0.0 to 1.0 for all combinations of models, prompts, and exams, except for GPT-3.5 using CoT prompting on the 100-question exam. As a result, the temperature hyperparameter of other LLMs may operate differently at temperatures above 1.0.
@@ 4.3 Implications
This research study provides empirical evidence that changes in sampling temperature in the range of 0.0 to 1.0 do not impact the problem-solving capabilities of LLMs on MCQA problems.
@@
This research also provides broader insights for AI researchers studying model hallucination and problem-solution state-space search with LLMs. Our results show that increasing LLM temperature up to 1.0 does not cause the LLM to hallucinate in ways that lead to incorrect MCQA solutions. In addition, higher temperatures do not appear to improve MCQA solution-space search in ways that lead to correct solutions more often.
@@ Appendix A Appendix
  C: Des Moines is a city located in the State of Iowa, but it is not the capital of Iowa.

### v3

@@ 1.1 Background
In addition, several inference hyperparameters can be adjusted to modify the LLM’s output at runtime. For example, hyperparameters like sampling temperature, top-k sampling, repetition penalty, and maximum token length all affect the LLM’s output and performance [3, 4, 5].
Despite significant interest in LLMs and progress in LLM best practices, many open questions remain about optimal prompt-engineering techniques and inference hyperparameters for LLMs. To complicate matters, various local optima may exist for LLMs, prompt types, and problem domains [5].
@@ 1.3 Choosing a Sampling Temperature
Practical guidelines for choosing a sampling temperature for a specific task or problem domain are often vague or anecdotal. Prompt-engineering guides often provide hypothetical examples of optimal sampling temperatures for various tasks. However, they rarely cite any sources or provide empirical evidence. [^1] A few empirical studies exist that indicate sampling temperature does have an effect on LLM performance on some types of problem-solving tasks (e.g., code generation, engineering exams, etc.) [11, 14, 12].
As a result, the current state of choosing the optimal sampling temperature for specific problems is largely based on guesswork, gut instinct, non-systematic experimentation, and iterative refinement. [^2] For example, OpenAI’s GPT-3.5 API allowed users to set the sampling temperature from 0.0 to 1.0 with a default of 0.7. GPT-4’s API expanded this range from 0.0 to 2.0 with a default of 1.0. No explanation from OpenAI has been provided for these default values or their change from GPT-3.5 to GPT-4 [15]., [^3] Even the GPT-4 Technical Report explains that the authors used their “best-guess” when choosing sampling temperatures while evaluating GPT-4 on various benchmarks. See Appendix A in the GPT-4 Technical Report [16].
@@ 2.1 Models
The models used in this research project comprise nine widely-used foundational LLMs. To complement our analysis, we also conducted experiments using five prompts created using commonly used prompt-engineering techniques.
@@ 2.3 Process
First, we instructed GPT-3.5 to complete the 100-question (small) exam using the CoT prompt with temperatures ranging from 0.0 to 2.0 in increments of 0.1. This allowed us to determine the range of viable sampling temperatures to explore.  [^5] For this experiment, we fixed all other sampling parameters (e.g., top-k, top-p, etc.) to isolate the effect of temperature.
@@ 3 Results
Llama 2 7B did not perform better than statistically random guesses. Its poor performance was due to generating incorrectly formatted answers (39%) and correctly formatted but incorrect answers (36%). Its all-or-nothing behavior at a temperature of 0.0 versus more random behavior from 0.1 to 1.0 led to a much lower, yet still non-significant, p-value.
@@
Finally, we analyzed the performance of GPT-3.5 using the CoT prompt on all ten exams. Accuracy remained stable for all temperatures across all problem domains based on visual analysis. However, the LSAT-AR and SAT-Math exams showed statistically significant differences in the Kruskal-Wallis p-values.  [^6] We considered the ARC Challenge results to be non-significant since they were greater than the significance threshold of 0.05. See Figure 5 and Table 5.
@@
We performed the Dunn-Bonferroni test on the LSAT-AR and SAT-Math results [49]. It revealed that the all-or-nothing behavior of responses generated at a temperature of 0.0 versus the more random responses from 0.1 to 1.0 caused the anomaly. The correct-answer accuracy for each individual problem varied significantly when evaluated pairwise across temperatures. However, the average accuracy for all problems remained similar across temperatures.
@@ 4.1 Interpretation
Based on these results, changes in temperature from 0.0 to 1.0 do not have a statistically significant effect on the problem-solving performance of LLMs. These results appear to generalize across LLMs, prompt-engineering techniques, and problem domains. However, there are a few potential exceptions to these general findings.
Therefore, we recommend setting an LLM’s sampling temperature to 0.0 for problem-solving tasks. This temperature maximizes reproducibility without compromising accuracy. In addition, it helps avoid the inevitable drop-off in performance that occurs beyond a temperature of 1.0. However, exceptions to this guidance should be taken into consideration.
@@ 4.2 Limitations
Fourth, due to time and cost constraints, we limited our study to two test sets of 1,000 and 100 randomly selected questions from standard benchmarks. These limited sample sizes may have introduced bias into the results. Utilizing a larger and more diverse test set would enhance the statistical reliability of our findings.
Fifth, we had to limit the sampling temperature range we explored from 0.0 to 1.0 for all combinations of models, prompts, and exams, except for GPT-3.5 using CoT prompting on the 100-question exam. As a result, the temperature hyperparameter of other LLMs may operate differently at temperatures above 1.0.
Sixth, we fixed all other sampling parameters (e.g., top-p, top-k, repetition penalty, etc.) to isolate the effect of sampling temperature. As a result, there may be combinations of sampling parameters that result in different outcomes.
@@ 4.3 Implications
This research study provides empirical evidence that changes in sampling temperature in the range of 0.0 to 1.0 do not significantly impact the problem-solving capabilities of LLMs on MCQA problems.
@@
This research also provides broader insights for AI researchers studying model hallucination and problem-solution state-space search with LLMs. Our results show that increasing LLM temperature up to 1.0 does not cause the LLM to hallucinate in ways that lead to incorrect MCQA solutions. In addition, higher temperatures do not appear to improve MCQA solution-space search in ways that lead to correct solutions more often than lower temperatures.
@@ Appendix A Appendix
  C: Des Moines is a city located in the State of Iowa and is the capital of Iowa.